OpenStudy (anonymous):

A cell phone company orders 500 new phones from a manufacturer. If the probability of a phone being defective is 2.6%, predict how many of the phones are likely to be defective. Round to the nearest whole number. A. 16 phones B. 130 phones C. 13 phones D. 11 phones

OpenStudy (anonymous):

2.6% of the phones are deffective, can you work out 2.6 of 500?

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

To work out a percentage of a number you do \[number \times \frac{ percentage }{ 100 }\]

OpenStudy (anonymous):

13 phones.

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

What about this one, it's the same just without a percentage.

OpenStudy (anonymous):

In a batch of 890 calculators, 12 were found to be defective. What is the probability that a calculator chosen at random will be defective? Write the probability as a percent. Round to the nearest tenth of a percent if necessary.

OpenStudy (anonymous):

A. 98.7% B. 1.3% C. 74% D. 1.8%

OpenStudy (anonymous):

So 12 out of 890 are defective. 12/890

OpenStudy (anonymous):

I got 0.013 in my calc.

OpenStudy (anonymous):

So 13?

OpenStudy (anonymous):

Wait. whoops. looking at the first questions answers lmfao

OpenStudy (anonymous):

so, 1.3?

OpenStudy (anonymous):

0.013 is the probability that you will get a defective one. To turn this into a percentage you multiply this by 100.

OpenStudy (anonymous):

Yep. I got 1.3 :D

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

I get this now. But can I do one more ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You have the numbers 1-24 written on slips of paper. if you choose one slip at random, what is the probability that you will not select a number which is divisible by 3? A. 2/3 B. 3/8 C. 1/3 D. 5/8

OpenStudy (anonymous):

Which numbers between 1 and 24 are divisible by 3? Well 3,6,9,12,15,18,21,24. So any of these 8 numbers out of the 24 will do the job

OpenStudy (anonymous):

So we could use 24 then?

OpenStudy (anonymous):

3x8 = 24

OpenStudy (anonymous):

So b, 3/8?

OpenStudy (anonymous):

Was wrong. :p

OpenStudy (anonymous):

No actually, 8 out of 24 is not 3/8

OpenStudy (anonymous):

Yeah that was my fault.

OpenStudy (texaschic101):

8/24 = 1/3